OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)]

OpenStudy (anonymous):

g(-2)= 6(-2)-1 g(-2)= -12-1 =-13 F(-13)= -2(-13)-1 26-1 =25

OpenStudy (anonymous):

thanks so much ive been having trouble with these equations

OpenStudy (anonymous):

It's alright, if you get another, solve for the g(x) first if possible, then take that answer and plug it in for the x in f(x)